# Double bracket vector fields on Poisson manifolds

### v1

@@ Abstract.
We generalize the double bracket vector fields defined on compact semi-simple Lie algebras to the case of general Poisson manifolds endowed with a pseudo-Riemannian metric. We construct a generalization of the normal metric such that the above vector fields, when restricted to a symplectic leaf, become gradient vector fields. We illustrate the discussion at a variety of examples and carefully discuss complications that arise when the pseudo-Riemannian metric does not induce a non-degenerate metric on parts of the symplectic leaves.
@@ 1. Introduction
Having a Riemannian manifold $(S,g^{S})$ and a cost function ${\widetilde{G}}\colon S\rightarrow\mathbb{R}$, one of the questions is, how can we compute the gradient vector field $\nabla_{{g}^{S}}{\widetilde{G}}$? To answer this, one needs to have sufficient information on the metric in order to solve the equation that defines the gradient vector field:
${g}^{S}(\nabla_{{g}^{S}}{\widetilde{G}},X)=\mathrm{d}{\widetilde{G}}(X),\quad\forall X\in\mathfrak{X}(S).$
This is oftentimes not possible or leads to very tedious computations, like in the case when one wishes to design steepest descend algorithms on manifolds.
A more amenable case is when $S$ is generalized in an ambient Riemannian manifold $(M,{g})$ with a simpler Riemannian geometry and the metric ${g}^{S}$ is the induced metric on $S$ by the ambient metric ${g}$. Knowing the way in which $S$ embeds into $M$, one can explicitly construct the orthogonal projection ${\textnormal{pr}}_{s}\colon T_{s}M\rightarrow T_{s}S$. Then
$\nabla_{{g}_{\textnormal{ind}}^{S}}{\widetilde{G}(s)}={\textnormal{pr}}_{s}~{}\nabla_{{g}}G(s),$
for all $s\in S$ and any smooth extension $G\colon M\rightarrow\mathbb{R}$ of ${\widetilde{G}}$. This approach has been extensively applied in optimization problems, e.g. in [1][12].
Another case is when we don’t know how the submanifold $S$ is generalized in the ambient space, but it can be described as a constraint submanifold. More precisely, let $S\subset M$ be a submanifold of a Riemannian manifold $(M,{g})$, which can be described by a set of constraint functions, i.e. $S={F}^{-1}(c_{0})$, where ${F}=(F_{1},\dots,F_{k})\colon M\rightarrow\mathbb{R}^{k}$ is a smooth map and $c_{0}\in\mathbb{R}^{k}$ is a regular value of ${F}$. As before, the submanifold $S$ becomes a Riemannian manifold when endowed with the induced metric $g_{\textnormal{ind}}^{S}$. We recall the so called embedded gradient vector field method to compute the gradient vector field of a cost function on $S$, using only the geometry of the ambient Riemannian manifold $M$ and the expressions of the constraint functions that describe the submanifold $S$.
Let $G\colon(M,{g})\rightarrow\mathbb{R}$ be a smooth extension of the cost function $\widetilde{G}$ on $S$ i.e. $\widetilde{G}=G|_{{F}^{-1}(c)}$. In [5], the following symmetric contravariant 2-tensor has been introduced:
$\mathbf{T}:=\sum_{i,j=1}^{k}(-1)^{i+j+1}\det\left(\text{Gram}_{(F_{1},...,\hat{F_{i}},...,F_{k})}^{(F_{1},...,\hat{F_{j}},...,F_{k})}\right)\nabla F_{i}\otimes\nabla F_{j}+\det\left(\text{Gram}_{(F_{1},...,F_{k})}^{(F_{1},...,F_{k})}\right){g}^{-1}.$ (1.1)
On the regular part of the foliation given by $F$, the generalized gradient vector field has been defined as follows:
$\partial_{g}G:=\frac{1}{\det\left(\text{Gram}_{(F_{1},\ldots,F_{k})}^{(F_{1},\ldots,F_{k})}\right)}i_{\mathrm{d}G}{\bf T},$ (1.2)
where
$\text{Gram}_{(g_{1},...,g_{r})}^{(h_{1},...,h_{r})}:=\left[\begin{array}[]{cccc}{g}(\nabla_{g}h_{1},\nabla_{g}f_{1})&...&{g}(\nabla_{g}h_{r},\nabla_{g}f_{1})\\ \vdots&\ddots&\vdots\\ {g}(\nabla_{g}h_{1},\nabla_{g}f_{r})&...&{g}(\nabla_{g}h_{r},\nabla_{g}f_{r})\end{array}\right].$
The generalized gradient vector field can be rewritten in the following form:
$\partial_{g}G(s)=\nabla_{g}G(s)-\sum\limits_{i=1}^{k}\sigma_{g}^{i}(s)\nabla_{g}F_{i}(s),\,\,\,\forall\,s\in M^{\textnormal{reg}},$
where the Lagrange multiplier functions $\sigma_{g}^{i}\colon M^{\textnormal{reg}}\rightarrow\mathbb{R}$ are defined by the formula:
$\sigma^{i}_{g}(s):=\frac{\det\left(\text{Gram}_{(F_{1},\ldots,F_{i-1},G,F_{i+1},\dots,F_{k})}^{(F_{1},\ldots,F_{i-1},F_{i},F_{i+1},...,F_{k})}(s)\right)}{\det\left(\text{Gram}_{(F_{1},\ldots,F_{k})}^{(F_{1},\ldots,F_{k})}(s)\right)}.$
As mentioned in [6], at a critical point $s_{0}\in S$ of $\widetilde{G}$, the numbers $\sigma^{i}_{g}(s_{0})$ coincide with the classical Lagrange multipliers.
Theorem 1.1 ( [5], [6]).
The generalized gradient vector field, defined in (1.2), is tangent to the regular part of the foliation given by $F$. More precisely, its restriction to a regular leaf is the gradient vector field
$\nabla_{g_{\textnormal{ind}}^{S}}{G}|_{{F}^{-1}(c)}(s)=\partial_{g}G(s),\,\,\,\forall s\in{F}^{-1}(c).$ (1.3)
Moreover, it is the unique vector field satisfying the above properties.
@@
Often, we don’t explicitly know enough components of the constraint map $F$ that defines the submanifold $S$. Such an instance appears when $S$ is a symplectic leaf of a Poisson manifold $M$ and we don’t know enough Casimirs to describe $S$. In the case of a linear Poisson structure, the gradient vector field on the symplectic leaf can be obtained as the restriction of a vector field on the ambient space $M$, namely the double bracket vector field. Unlike in the previous setting, the metric on the symplectic leaf will not be the induced one, but the so called normal metric.
@@
We recall the classical case of double bracket vector field on a semi-simple Lie algebra. Let $(\mathfrak{g},[\ ,\ ])$ be a semi-simple Lie algebra with $\kappa\colon{\mathfrak{g}}\times{\mathfrak{g}}\to\mathbb{R}$ the Killing form, hence a non-degenerate, symmetric, Ad-invariant bilinear form. The following vector field, called the double bracket vector field, has been introduced by Brockett [10], [11], see also [7], in the context of dynamical numerical algorithms and linear programming:
@@
Let $N$ be a fixed regular element in the compact semi-simple Lie algebra $\mathfrak{g}$. For the smooth function $H\colon\mathfrak{g}\to\mathbb{R}$ defined by $H(L)={\kappa}(L,N)$ the following equality holds:
@@
Our goal is to obtain an analogue of the above theorem in the setting of a general Poisson manifold $(M,\Pi)$ equipped with a Riemannian or pseudo-Riemannian metric $g$.
The structure of the paper is as follows: In Sec. 2 we construct generalized double bracket vector fields for an arbitrary triple $(M,g,\Pi)$. These vector fields are tangent to the symplectic leaves of $(M,\Pi)$ and are analogous to the double bracket vector fields (1.4) in the linear case of a semi-simple Lie algebra $\mathfrak{g}$; they are reduced to those for $M=\mathfrak{g}^{*}$, equipped with its canonical Poisson structure $\Pi_{lin}$ and the Killing metric. Following an idea presented in [5], we construct a metriplectic tensor that couples the Poisson structure and the pseudo-Riemannian structure, analogous to (1.1).
In Section 3, we construct a generalized normal pseudo-Riemannian metric, that we call double bracket metric. Its construction is limited to “good leaves” of “green zones”—notions we define in the body of the paper. We then show that the restriction of the generalized double bracket vector field to a good symplectic leaf or a green zone in a symplectic leaf is the gradient of a smooth function with respect to this double bracket metric.
The Section 4 contains a large class of examples that illustrate all the new notions introduced in the paper and their subtle differences.
In Subsection 4.1, we provide a large class of Poisson structures on $\mathbb{R}^{3}$ and study their symplectic leaves. These include the Poisson-Lie manifold $\mathfrak{sl}_{2}^{*}$ as well as a related Poisson-Lie group. But they are much more general and, e.g., permit one to construct examples of symplectic leaves of arbitrary genus, as explained in Subsection 4.2.
In Subsection 4.3, we introduce the pseudo-metric $g$ which we put on the Poisson manifolds of the preceding section. This then permits us to discuss where problems with the non-degeneracy of $g$ pulled back to symplectic leaves arise. We call the set $R$ of such points the “red zone”, as it prohibits applying the constructions of Section 3.
In Subsection 4.4, the information about symplectic leaves gathered in Subsection 4.1 is combined with the one on the red zone. Intersecting a leaf $S$ with $R$ yields what we call “red lines”; the remainder of the leaf, $S\backslash(R\cap S)$, then provides the “green zones”. Leaves not having any such an intersection are simply “good leaves”. It is precisely the good leaves or, more generally, the green zones, where our main theorem will be applicable.
Finally, in Subsection 4.5, we compute the structures induced on the leaves for the class of our examples and illustrate the main theorem by means of them.
2. Generalized double bracket vector field
Let $(M,{g})$ be a pseudo-Riemannian manifold that is also endowed with a Poisson bracket
{,}$\{\cdot,\cdot\}\colon{\cal C}^{\infty}(M)\times{\cal C}^{\infty}(M)\to{\cal C}^{\infty}(M).$
@@ 2. Generalized double bracket vector field
We call the metriplectic tensor the following symmetric contravariant 2-tensor ${\mathcal{M}}\colon\Omega^{1}(M)\times\Omega^{1}(M)\to{\cal C}^{\infty}(M)$,
$\mathcal{M}(\alpha,\beta):={g}(\sharp_{{}_{\Pi}}\alpha,\sharp_{{}_{\Pi}}\beta)\,,$
[…]
@@
Since $\nabla G(L)=\mathrm{d}\bar{G}(\xi)$, where $\mathrm{d}\bar{G}(\xi)\in\mathfrak{g}^{**}=\mathfrak{g}$, the identity (2.4) follows from (2.5) and (2.6).
@@ 3. Gradient nature of the generalized double bracket vector field
Let $(M,\Pi,g)$ be a Poisson manifold endowed with a pseudo-Riemannian metric. In this section, we prove that the generalized double bracket vetor field defined in Definition 2.2, tangent to the symplectic leaves, is a gradient vector field with respect to a metric, that we call double bracket metric. It generalizes the normal metric defined on the regular adjoint orbits of a semi-simple compact Lie algebra. In the case when the ambient metric is of indefinite signature, one has to be careful when dealing with submanifolds as they might not be pseudo-Riemannian with respect to the induced metric [17].
@@
Let $S$ denote a symplectic leaf of $(M,\Pi,g)$ and $\iota\colon S\hookrightarrow M$. Then the two-tensor $g_{\textnormal{ind}}^{S}:=\iota^{*}g$ induced by $g$ on $S$ is degenerate at $s\in S$ if and only if $s$ is $\mathcal{M}$-singular.
@@
In what follows we will define the double bracket metric on a good symplectic leaf and we will prove that this metric generalizes the normal/standard metric on adjoint orbits of a semi-simple compact Lie algebra. However, all the statements below will be applicable also to $\mathcal{M}$-regular regions of a symplectic leaf $S$, called green zones in Section 4.4 below.
@@
4. A case study
In this section we work with a large class of Poisson structures on $\mathbb{R}^{3}$ that illustrate all the new notions introduced in the previous sections. Among them, we specify three examples of a linear Poisson structure, a quadratic Poisson structure and a Poisson Lie group. Symplectic leaves and green zones, where our main theorem is applicable, in each case is carefully studied.
4.1. A class of Poisson structures on $\mathbb{R}^{3}$ and their symplectic leaves
The canonical basis of the Lie algebra $\mathfrak{sl}_{2}$,
@@ 4.1. A class of Poisson structures on $\mathbb{R}^{3}$ and their symplectic leaves
Figure 1. Symplectic leaves on $\mathfrak{sl}(2)^{*}$: (A) for $c=-1$, (B) for $c=0$, and (C) for $c=1$
@@
4.2. Topological nature of certain symplectic leaves
@@
Figure 5. (A) Function $h_{c}$ for the linear Poisson structure on $\mathfrak{sl}^{*}_{2}$, orange color for $h_{-1}$, black color for $h_{0}$ and blue color for $h_{1}$. (B) Function $h_{c}$ for the quadratic Poisson structure, orange color for $h_{0}$ and blue color for $h_{1}$.
We see that $h_{-1}$ has no zeros and that the corresponding leaf is topologically trivial, cf. Fig. 1A, while $h_{1}$ has two simple roots and the leaf $S$ for $c=1$ is a cylinder topologically. This qualitative behavior of $h_{c}$ and the corresponding leaves remains if one does not change the sign of $c$. For the special value $c=0$, on the other hand, we have one multiple root of this function and precisely there we find a singular leaf together with two regular ones.
@@
4.3. Degeneracy structure of the metriplectic tensor $\mathcal{M}$: red zones
The second main ingredient is the choice of a metric on the space where the leaves are embedded in, i.e. here on the Poisson manifold $\mathbb{R}^{3}$. In the case of the linear $\mathfrak{sl}_{2}^{*}$ brackets, there is a natural candidate for such a metric, the Killing metric $\kappa$ of the Lie algebra $\mathfrak{sl}_{2}$. It is of indefinite signature, corresponding to the fact that $\operatorname{SL}(2,\mathbb{R})$ is a non-compact Lie group. After a trivial rescaling, $g:=\tfrac{1}{2}\kappa$, this metric takes the form

### v2

# Generalized double bracket vector fields

@@ Abstract.
By introducing a metriplectic tensor, we generalize the double bracket vector fields defined on semi-simple Lie algebras to the case of Poisson manifolds endowed with a pseudo-Riemannian metric. We also construct a generalization of the normal metric on an adjoint orbit of a compact semi-simple Lie algebra such that the above vector fields, when restricted to a symplectic leaf, become gradient vector fields. We illustrate the newly introduced objects to a variety of examples and carefully discuss complications that arise when the pseudo-Riemannian metric does not induce a non-degenerate metric on parts of the symplectic leaves.
@@ 1. Introduction
On a Riemannian manifold, one of the questions is how can one compute the gradient vector field of a cost function. This problem appears, for instance, in designing steepest descent algorithms on Riemannian manifolds, like matrix Lie groups and homogeneous spaces. To answer this, one needs to have sufficient information on the metric in order to solve the equation that defines the gradient vector field. This is oftentimes not possible or leads to very tedious computations.
A more amenable case is when our manifold of interest is realized as a submanifold in an ambient Riemannian manifold with a simpler Riemannian geometry, so that one can explicitly construct the orthogonal projection operator that relates the gradient vector fields in the two manifolds. This approach has been extensively applied in constrained optimization problems, e.g. in [1][13].
In the often encountered case of a submanifold described by a set of constraint functions, the gradient vector field on the submanifold, with respect to the induced metric, can be realized as a restriction of a vector field (called embedded gradient vector field) defined on the ambient space. This construction has been worked out in [5][6].
@@
Sometimes it happens that not enough components of the constraint map that defines the submanifold are available. This case appears, for instance, when one does not explicitly know enough Casimirs to describe a symplectic leaf of a Poisson manifold. Nevertheless, in the case of the Lie-Poisson structure for a compact Lie algebra, the gradient vector field on the regular symplectic leaves has been obtained as the restriction of a vector field on the ambient space, namely the double bracket vector field. Unlike in the previous setting, the metric on the symplectic leaf will not be the induced one, but the so called normal metric.
Our goal is to obtain an analogue of the above scenario in the setting of a general Poisson manifold equipped with a (pseudo-)Riemannian metric.
More precisely, following an idea presented in [5], we construct a symmetric contravariant tensor that couples the Poisson structure and the Riemannian structure (no compatibility conditions required). This type of tensors are called in the literature metriplectic tensors [16][17][20]. Our newly constructed tensor allows to associate a vector field to every smooth function in a naturally way. We call this vector field the generalized double bracket vector field as, for the case of a compact semi-simple Lie algebra, we recover the classical double bracket vector field. The generalized double bracket vector field has some useful properties. It is tangent to all the symplectic leaves. Moreover, when restricted to such a leaf, it proves to be of gradient type with respect to a natural metric that generalizes the normal metric.
All the constructions above work more generally for a pseudo-Riemannian structure on the Poisson manifold. Thus, non-compact semi-simple Lie algebras can be included. The trade off is that in this case one needs to restrict to what we will call good symplectic leaves, namely those for which the induced metric is non-degenerate.
However, new complications arise in the nonlinear Poisson setting. An important question is if the restriction of the indefinite signature metric to the symplectic leaves is non-degenerate. While for the special case of a non-compact semi-simple Lie algebra this happens on some exceptional leaves only, in the general case, the situation is more intricate and this question will occupy us for a good part of this paper. We will in particular look at the case of the Lie algebra $\mathfrak{sl}(2,\mathbb{R})$ and a huge class of Poisson structures on $\mathbb{R}^{3}$ generalizing it non-linearly, while keeping a rescaling of the Killing form of $\mathfrak{sl}(2,\mathbb{R})$ in the ambient space $\mathbb{R}^{3}$. This will lead to an interesting and sophisticated interplay between the Poisson and the pseudo-Riemannian structures.
Our constructions that generalize the double bracket vector fields are suitable for the study of various dynamical properties of dissipative systems, analogous to other methods, e.g. those used in [7][18].
Structure of the paper
We start Section 2 by recalling the classical double bracket vector field in the linear case of a semi-simple Lie algebra $\mathfrak{g}$. In order to generalize this setting, following an idea presented in [5], we construct a metriplectic tensor that couples a Poisson structure with a pseudo-Riemannian structure. We introduce the generalized double bracket vector field and prove that in the particular case of a semi-simple Lie algebra it becomes the classical double bracket vector field.
In Section 3, we construct a pseudo-Riemannian metric on symplectic leaves, that we call double bracket metric. This generalizes the normal metric defined on an adjoint orbit of a compact semi-simple Lie algebra. Its construction is limited to good symplectic leaves, or more generally to green zones of symplectic leaves, notions that we define in the body of the paper. We then show that the restriction of the generalized double bracket vector field to a good symplectic leaf or a green zone in a symplectic leaf is the gradient of a smooth function with respect to this double bracket metric.
In Section 4, we provide a large class of Poisson structures on $\mathbb{R}^{3}$ and study their symplectic leaves. These include the Lie-Poisson manifold $\mathfrak{sl}(2,\mathbb{R})^{*}$ as well as a related Poisson-Lie group. But they are much more general and, e.g., permit one to construct examples of symplectic leaves of arbitrary genus as explained in Section 5.
In Section 6, we introduce the pseudo-metric $g$ which we put on the Poisson manifolds of the preceding section. This then permits us to discuss where problems with the non-degeneracy of $g$ pulled back to symplectic leaves arise. We call the set $R$ of such points the red zone, as it prohibits applying the constructions of Section 3.
In Section 7, the information about symplectic leaves gathered in Section 4 is combined with the one on the red zone. Intersecting a leaf $S$ with $R$ yields what we call red lines; the remainder of the leaf, $S\backslash(R\cap S)$, then provides the green zones. Leaves not having any such an intersection are simply good leaves. It is precisely the good leaves or, more generally, the green zones, where the Theorem 3.4 will be applicable. In the final Section 8 we compute the structures induced on the leaves for the class of our examples and illustrate Theorem 3.4 by means of them.
2. Metriplectic tensor. Generalized double bracket vector field
In this section we introduce the metriplectic tensor and the generalized double bracket vector field. These objects are constructed in a very general setting of a Poisson manifold $(M,\Pi)$ endowed with a pseudo-Riemannian metric $g$, with no compatibility conditions required. We recover, as a particular case, the classical double bracket vector field.
@@
We start by recalling the classical setting of the double bracket vector field. Let $(\mathfrak{g},[\ ,\ ])$ be a semi-simple Lie algebra with $\kappa\colon{\mathfrak{g}}\times{\mathfrak{g}}\to\mathbb{R}$ the Killing form, hence a non-degenerate, symmetric, Ad-invariant bilinear form. The following vector field, called the double bracket vector field, has been introduced by Brockett [11], [12], see also [8], in the context of dynamical numerical algorithms and linear programming:
@@
Let $N$ be a fixed regular element in the compact semi-simple Lie algebra $\mathfrak{g}$. The gradient of the linear function $H\colon\mathfrak{g}\to\mathbb{R}$ defined by $H(L)={\kappa}(L,N)$ restricted to a regular adjoint orbit $S$, taken with respect to the normal metric, is
@@
We move to a more general case: $(M,\Pi,{g})$ a Poisson manifold endowed with a pseudo-Riemannian metric.
@@
We call metriplectic tensor the following symmetric contravariant 2-tensor ${\mathcal{M}}\colon\Omega^{1}(M)\times\Omega^{1}(M)\to{\cal C}^{\infty}(M)$,
$\mathcal{M}(\alpha,\beta):={g}(\sharp_{{}_{\Pi}}\alpha,\sharp_{{}_{\Pi}}\beta)\,.$
@@
Since $\nabla G(L)=\mathrm{d}\bar{G}(\xi)$, where $\mathrm{d}\bar{G}(\xi)\in\mathfrak{g}^{**}=\mathfrak{g}$, the identity (2.7) follows from (2.8) and (2.9). ∎
@@ 3. Gradient nature of the generalized double bracket vector field
In this section, we prove that the generalized double bracket vector field defined in Definition 2.2, when restricted to a leaf, is a gradient vector field with respect to a metric, that we call double bracket metric. This metric generalizes the normal metric defined on the regular adjoint orbits of a semi-simple compact Lie algebra. In the case when the ambient metric is of indefinite signature, one has to be careful when dealing with submanifolds as they might not be pseudo-Riemannian with respect to the induced metric [19].
@@
Let $S$ denote a symplectic leaf of $(M,g,\Pi)$ and $\iota\colon S\hookrightarrow M$. Then the two-tensor $g_{\textnormal{ind}}^{S}:=\iota^{*}g$ induced by $g$ on $S$ is degenerate at $s\in S$ if and only if $s$ is $\mathcal{M}$-singular.
@@
In what follows we will define the double bracket metric on a good symplectic leaf and we will prove that this metric generalizes the normal/standard metric on adjoint orbits of a semi-simple compact Lie algebra. However, all the statements below will be applicable also to $\mathcal{M}$-regular regions of a symplectic leaf $S$, called green zones in Section 6 below.
@@
4. A class of Poisson structures on $\mathbb{R}^{3}$ and their symplectic leaves
In this section we work with a large class of Poisson structures on $\mathbb{R}^{3}$ that illustrate all the new notions introduced in the previous sections. Among them, we specify three examples of a linear Poisson structure, a quadratic Poisson structure and a Poisson Lie group. Symplectic leaves and green zones of each case, where our main theorem is applicable, are carefully studied in the next sections.
The canonical basis of the Lie algebra $\mathfrak{sl}(2,\mathbb{R})$,
@@ 4. A class of Poisson structures on $\mathbb{R}^{3}$ and their symplectic leaves
Figure 1. Symplectic leaves on $\mathfrak{sl}(2,\mathbb{R})^{*}$: (A) for $c=-1$, (B) for $c=0$, and (C) for $c=1$
@@
5. Topological nature of certain symplectic leaves
@@
Figure 5. (A) Function $h_{c}$ for the linear Poisson structure on $\mathfrak{sl}(2,\mathbb{R})^{*}$, orange color for $h_{-1}$, black color for $h_{0}$ and blue color for $h_{1}$. (B) Function $h_{c}$ for the quadratic Poisson structure, orange color for $h_{0}$ and blue color for $h_{1}$.
We see that $h_{-1}$ has no zeros and that the corresponding leaf is topologically trivial, cf. Fig. 1A, while $h_{1}$ has two simple roots and the leaf $S$ for $c=1$ is a cylinder topologically. This qualitative behaviour of $h_{c}$ and the corresponding leaves remains if one does not change the sign of $c$. For the special value $c=0$, on the other hand, we have one multiple root of this function and precisely there we find a singular leaf together with two regular ones.
@@
6. Degeneracy structure of the metriplectic tensor: red zones
The second main ingredient is the choice of a metric on the space where the leaves are embedded in, i.e. here on the Poisson manifold $\mathbb{R}^{3}$. In the case of the linear $\mathfrak{sl}(2,\mathbb{R})^{*}$ brackets, there is a natural candidate for such a metric, the Killing metric $\kappa$ of the Lie algebra $\mathfrak{sl}(2,\mathbb{R})$. It is of indefinite signature, corresponding to the fact that $\operatorname{SL}(2,\mathbb{R})$ is a non-compact Lie group. After a trivial rescaling, $g:=\tfrac{1}{2}\kappa$, this metric takes the form